OpenStudy (anonymous):

Eight less than a number s is three more than four times a number t. Solver for s

jhonyy9 (jhonyy9):

s>8 3>4t t<3/4

jhonyy9 (jhonyy9):

you have tought in this way ?

OpenStudy (anonymous):

s-8=3+4t s= 4t+11